OpenStudy (anonymous):

[2.03] Solve: −4(x + 1) − 3 = −3(x − 4) −19 19 2 0

OpenStudy (anonymous):

[2.03] ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

-4 x-7 = 12-3 x

OpenStudy (anonymous):

First you should distribute -4(x+1) and -3(x-4) Remember how to do that?

OpenStudy (anonymous):

x=-19

OpenStudy (anonymous):

@MathSofiya yes?

OpenStudy (anonymous):

its steps -4 (x+1)-3 = 12-3 x

OpenStudy (anonymous):

can you please help me when you're done. I'm really sorry to interrupt.

OpenStudy (anonymous):

@best.shakir It's better if we don't give the answers. We want the student to learn, not just to plug the correct answer on the HW.

OpenStudy (anonymous):

i gave the steps also

OpenStudy (anonymous):

But you should ask him/her to try it before you do the whole problem.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@best.shakir Thanks! Your The Best!

OpenStudy (anonymous):

welcom....